OpenStudy (yrelhan4):

Optics question. I know all the formulas and stuff, but i dont know how to approach this one.

OpenStudy (yrelhan4):

http://i.imgur.com/TcW184K.png

OpenStudy (dumbcow):

sorry don't know this one either....sounds like you just plug it into a formula i dont know the formulas for optics?

OpenStudy (yrelhan4):

i'm confused with which formula to apply tho.. i have these two formulas [1-(1/mu)]t=shift and Happ=hac*mu or happ=hac/mu, for rays going from rarer to denser medium and from denser to rarer medium respectively.. i am confused lol.

OpenStudy (anonymous):

could you type out your formulas with the equation button? it would make them easier to identify vvv

OpenStudy (yrelhan4):

well i hate using the equation button.. i'll write the down.

OpenStudy (yrelhan4):